Voltage and current in an AC circuit are given by
V = 5 sin (wt – pi/6)
I = 4 sin (wt + pi/6)
How much volatge is leading?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
current lead the voltage by = pi/ 6 – (- pi/6) = pi/3
 
or we can say that voktage leads by – 30 degrees